# OpenRubrics: Towards Scalable Synthetic Rubric Generation for Reward Modeling and LLM Alignment

### v1

@@ 5.1 Datasets and Experiment Settings
We train both components of Rubric-RM—the rubric generator and the judge—on the curated OpenRubrics as presented in Sec. 4.2. Rubrics are produced with contrastive signals from chosen/rejected responses and filtered by preference–label consistency before use. Unless otherwise noted, we use the science-related slice of OpenRubrics to better match our domain study on HealthBench/medical evaluation.
Backbone and variants.
@@
Both the rubric generator and the judge are fine-tuned from Qwen-3-8B (“Rubric-RM-8B”) unless specified. At inference time, Rubric-RM follows a two-stage pipeline: (i) generate or retrieve a rubric conditioned on the prompt and candidate responses; (ii) predict the pairwise preference conditioned on that rubric. We also report an ensemble variant, voting@5, which aggregates five independently sampled judge trajectories by majority vote.
Baselines.
@@
We compare against strong, same-scale white-box reward/judge models—JudgeLRM-7B (Chen et al., 2025a), RRM-7B (Guo et al., 2025b), and RM-R1-7B (Chen et al., 2025b)—as well as larger RM-R1-14B (Chen et al., 2025b) and reference API judges when available. To isolate the benefit of rubric-aware fine-tuning, we also include a naïve pipeline Qwen-3-8B (Rubric+Judge) that directly prompts the base model to produce a rubric and then make a judgment.
Evaluation benchmarks and metrics.
@@
We evaluate Rubric-RM as a pairwise reward model on widely used reward-modeling suites: RewardBench (Chat / Chat-Hard) (Lambert et al., 2025b), RM-Bench (Liu et al., 2025c), PPE-IFEval (Frick et al., 2024b), FollowBench (Jiang et al., 2024), InfoBench (Qin et al., 2024), IFBench (Peng et al., 2025), and RewardBench2 (Precise-IF / Focus) (Malik et al., 2025). While FollowBench and InfoBench were originally designed to assess instruction-following capabilities of LLMs, we adapt them into pairwise evaluation settings by sampling two responses from the same model (Qwen-3-8B/14B), where one response adheres to all specified constraints and the other violates some of them. For the domain study we additionally report HealthBench/medical results. We follow each benchmark’s official splits and scoring rules, reporting accuracy/win-rate or benchmark-defined scores.
Decoding and efficiency protocol.
All models are run under matched decoding budgets (temperature, max tokens, and stop conditions per benchmark recommendations). We use unified execution stack vLLM (Kwon et al., 2023) for throughput-fair comparisons. For efficiency (Table 4), we measure wall-clock time to score a fixed set of prompts; note that rubrics from stage (i) are cacheable and can be reused across examples, amortizing the cost in large-scale judging and preference optimization.
Policy-model evaluation.
When integrating Rubric-RM into policy optimization, we follow prior work that evaluates on instruction-following (IF) suites (Zhou et al., 2023, Qin et al., 2024, Viswanathan et al., 2025), using each benchmark’s official metrics and scripts.
Reproducibility.
We use LLaMA-Factory (Zheng et al., 2024) for training Rubric-RM (via SFT) and policy models (via DPO; Rafailov et al. (2023)). For evaluation, we use the benchmarks’ official scripts where available. To facilitate reproducibility, we release our training and inference configuration in Appendix A.1. Prompts, including rubric templates, are provided in Appendix A.2.
5.2 Performance of Rubric-RM
We first validate the performance of Rubric-RM for reward modeling. For a more systematic evaluation, we test both 4B and 8B variants of Rubric-RM, which use Qwen3-4B and Qwen3-8B as backbones respectively. Table 1 reports the performance of our proposed Rubric-RM across multiple benchmarks.
Outperforming Comparable Reward Models.
Both Rubric-RM-4B and Rubric-RM-8B outperform existing 7B-scale white-box reward models such as JudgeLRM-7B, RRM-7B, and RM-R1-7B. Rubric-RM-4B achieves an average score of 65.6, already higher than JudgeLRM-7B (53.8), RRM-7B (57.8), and RM-R1-7B variants (59.4–61.7), while Rubric-RM-8B further improves to 68.5. These results demonstrate that rubric-aware training produces more reliable and generalizable reward signals even under smaller model scales, outperforming models trained with generic preference-based supervision.
Majority Voting Further Enhances Performance.
We also evaluate Rubric-RM-voting@5, which aggregates predictions via majority voting across five independent judge trajectories. This ensemble strategy consistently improves accuracy. Rubric-RM-4B-voting@5 reaches 68.3, and Rubric-RM-8B-voting@5 achieves the best overall average of 71.2, nearly matching much larger models such as RM-R1-14B (71.7) and the Rubric+Judge API (71.3). These results highlight the robustness and stability benefits of rubric-based ensembles.
Effectiveness of Rubric-Aware Fine-Tuning.
A direct pipeline of generating rubrics with Qwen-3-8B and then applying them for judgment performs poorly (average 58.9). In contrast, our Rubric-RM substantially outperforms this baseline, achieving 68.5 on average. This demonstrates that our fine-tuning with high-quality rubrics, generated by contrastive responses and filtered by preference–label consistency, provides a significant advantage over naïve rubric-based judging.
Strength on IF Evaluation Benchmarks.
In addition to absolute improvements, Rubric-RM shows particularly strong advantages on the IF Evaluation Benchmarks, which measure fine-grained instruction following capabilities. For example, on FollowBench and InfoBench, Rubric-RM achieves 73.1 and 78.6, respectively, substantially outperforming other 7B-scale baselines such as JudgeLRM-7B (79.8 / 62.7) and RRM-7B (65.5 / 68.2). These results demonstrate that rubric-based training is especially effective at capturing instruction adherence and nuanced response qualities, where conventional reward models often struggle.
In the remaining part of experimentation, we use Rubric-RM-8B as our reward model unless specified.
RewardBench IF Evaluation Benchmarks RM-Bench RewardBench2 HelpSteer3 AVG
@@ 5.2 Performance of Rubric-RM
Qwen-3-8B (Rubric+Judge) 73.9 63.6 63.0 53.8 74.6 55.6 64.2 21.9 56.6 61.8 58.9
@@
Rubric-RM-4B 83.2 67.4 71.1 61.8 77.4 62.1 61.1 28.1 76.6 66.7 65.6
Rubric-RM-4B-voting@5 83.6 69.1 77.3 65.6 80.9 64.4 61.8 34.1 78.2 68.2 68.3
Rubric-RM-8B 87.3 73.0 73.1 65.1 78.6 64.6 62.2 33.8 78.2 68.6 68.5
Rubric-RM-8B-voting@5 89.6 75.7 78.2 68.0 81.6 67.1 62.4 40.0 79.2 69.9 71.2
Table 2: Comparison of trained policy models with different reward models on a format-based constrained instruction-following benchmark (IFEval) and an open-ended benchmark (InfoBench). Baseline results are from Viswanathan et al. (2025). Results with underlines are reproduced by us using official checkpoints and evaluation scripts. Best scores are in bold.
@@
AutoIF (Dong et al., 2024) 56.9 47.1 67.0 57.6 57.2 80.6
UltraIF (An et al., 2025) 75.4 71.3 83.0 79.4 77.3 80.7
@@
+ DPO (via Rubric-RM) 79.1 73.8 85.3 81.3 79.9 82.9
Table 3: Comparison of different alignment strategies applied to Qwen2.5-7B-Instruct on Arena-Hard and AlpacaEval benchmarks. Results are reported for vanilla models and style/length-controlled settings. Baseline results are from Viswanathan et al. (2025). Best results are in bold.
@@
Vanilla Style-Controlled Vanilla Length-Controlled
@@
+ DPO (via Rubric-RM) 56.1 56.9 38.8 50.5 50.6
5.3 Offline Reinforcement Learning for Policy Models with Rubric-RM
@@ 5.3.1 Instruction-Following Evaluation
We further evaluate the effectiveness of using Rubric-RM as a reward model for policy optimization on instruction-following tasks, including IFEval, InfoBench, and IFBench. The results are shown in Table 2 and Fig. 4.
Improved Performance on IFEval and InfoBench.
When serving as the reward model in Direct Preference Optimization (DPO), Rubric-RM enables the trained policy model to achieve the best overall performance among all open-source counterparts. On IFEval, the Rubric-RM-trained policy achieves an average score of 79.9, surpassing policies trained with Skywork (76.0) and ArmoRM (76.0). On InfoBench, the Rubric-RM-based policy reaches 82.9, outperforming other DPO-trained policies and approaching the performance of much larger commercial systems. These results highlight that rubrics provide more reliable optimization signals for constrained instruction-following.
Clear Gains on Complex Instruction Following Benchmark (IFBench).
Figure 4 shows that the policy model optimized with Rubric-RM attains a new state-of-the-art score of 33.7 on IFBench, substantially higher than RLCF (28.2) and RLMT-based methods (22.4–22.8). Compared with both supervised fine-tuning variants and reinforcement learning baselines, Rubric-RM provides stronger inductive biases, enabling policies to better capture fine-grained instruction adherence.
Overall, these results confirm that using Rubric-RM as a reward model significantly enhances the instruction-following ability of trained policies. Unlike prior scalar or generative reward models that often fail to enforce strict formats or nuanced constraints, rubrics supply explicit and interpretable guidance, leading to consistent improvements across multiple benchmarks. This demonstrates that Rubric-RM-based training not only improves absolute performance but also establishes a robust foundation for building policy models aligned with human instructions.
Figure 4: Comparison of trained policy models on IFBench. Results of baselines except RLCF are from Bhaskar et al. (2025). We evaluate RLCF with its official checkpoint.
5.3.2 Human Preference Alignment Evaluation
We evaluate policies trained with Rubric-RM on human preference alignment benchmarks Arena-Hard and AlpacaEval (Table 3).
With DPO optimization, Rubric-RM achieves the best overall average score (50.6) among all open-source reward models. On Arena-Hard (style-controlled), it obtains 56.9, outperforming Skywork (50.3), Ultrafeedback (47.9), and RLCF (48.4). On AlpacaEval (length-controlled), it reaches 50.5, surpassing ArmoRM (38.1) and AI Judge (33.4). These results show that rubric-based signals provide reliable gains across both vanilla and controlled settings.
@@ 5.4 Rubric-RM for BioMedical Domain
In this section we further study the effectiveness of Rubric-RM in more specialized medical domain, following Arora et al. (2025). Experiments here are again conducted from two aspects: (i) how OpenRubrics results in better reward modeling ability in Rubric-RM; (ii) how does Rubric-RM leads to stronger policy models. The Rubric and Judge are Qwen-3-8B backbones fine-tuned with OpenRubrics data from science-related domains, see Sec 4.1 for more details about our data.
5.4.1 Performance of Rubric-RM on HealthBench
As in general domains, Rubric-RM outperforms comparable-size generative reasoning reward models on HealthBench: our model obtains 68.3, exceeding RRM-7B (63.3) and two 7B-level variants of RM-R1 (55.4/66.9). Notably, Rubric-RM matches the competitiveness of much larger RM-R1-14B (69.9). Moreover, consistent with previous results, majority voting further enhances Rubric-RM performance: Rubric-RM-voting@5 attains 72.9 (+4.6 over single-pass judge), narrowing the gap to much larger 14B reasoning models (e.g., RM-R1-14B up to 74.7) and approaching API-based references (69.9 – 73.5). These gains clearly reflect the effectiveness of Rubric-RM trained with OpenRubrics.
The second noticeable gain lies in the importance of domain-specific SFT. Compared to directly prompting Qwen-3-8B with the proposed “Rubric+Judge” pipeline, which only reaches score 51.8, Rubric-RM (68.3) yields remarkable improvement on HealthBench (+16.5). This pronounced gap highlights the importance of domain-specific rubric data and rubric-aware SFT: contrastive rubric training and preference–label consistency produce higher-precision, science-aware criteria that transfer more effectively to health tasks than rubrics generated on-the-fly.
@@ 5.4.1 Performance of Rubric-RM on HealthBench
5.4.2 Preference Optimization with Rubric-RM on HealthBench
We further validate that our Rubric-RM, which achieves higher reward modeling performance as presented in the previous section, can successfully translate into stronger policy model learning. Here we compare using Rubric-RM as the preference judge for DPO on HealthBench against two baselines expanding both reasoning-based RM-R1-7B (Qwen-2.5-Inst), and non-reasoning based ArmoRM (Wang et al., 2024a). Specifically, we use Qwen-2.5-7B-Instruct as the base policy model and collect independent 4 responses from it to each question in HealthBench. The preference pairs are labeled with different reward models, upon which the base model is finetuned with DPO.
The DPO performance results are reported in Figure 5(b). According to the Figure, holding the policy backbone and DPO recipe fixed, replacing baseline judges with Rubric-RM consistently yields the best downstream performance. Starting from the base model (21.6), DPO via ArmoRM reaches 22.5 and via RM-R1-7B reaches 22.7, whereas DPO via Rubric-RM attains 23.8, the highest among all settings. This results in 1.1 to 1.3 absolute gain over strong 7B reasoning rewards echoes our findings that rubric-aware, domain-tuned signals provide cleaner preferences for policy optimization than generative reasoning at a similar scale.
@@ 5.5 Efficiency Comparison
In this section we analyze the computing cost for Rubric-RM inference. Table 4 reports wall-clock time on 100 randomly sampled prompts from RewardBench2. Notably, our Rubric-RM (130.77 sec), whose rubric generator and judge are both Qwen-3-8B, is not slower than existing reasoning reward models such as RRM-7B (203.4 sec) and RM-R1-7B/14B (170.76 – 382.02 sec). In particular, it is markedly faster than the 14B R1 variants and competitive with the stronger 7B reasoning baselines, while operating at the 8B scale. We attribute the speed gap to the different computation patterns. Prior reasoning reward models execute long chain-of-thought traces before emitting a final judgment, incurring substantial decoding latency. In contrast, our approach decomposes evaluation into two focused stages—(i) generating or retrieving a rubric and (ii) applying a lightweight judge conditioned on that rubric—so each step remains short and targeted. A further practical advantage is that rubrics are amortizable: once produced, they can be computed offline and cached for reuse across many examples, eliminating rubric generation cost at scoring time. This property makes Rubric-RM especially attractive for large-scale preference optimization, where repeated judgments dominate runtime. While JudgeLRM-7B achieves the lowest raw latency (25.71 sec), it does not provide the explicit rubric signal that enables our method’s interpretability and downstream policy optimization.
5.6 Case Study
Case Study: Instruction adherence on RewardBench Chat Hard
Prompt Describe a vivid and unique character using strong imagery, in fewer than two paragraphs.
Resp B (snippet) “ …a fierce and determined young woman…[paragraph break] …Which of the following is an example of a character…[paragraph break]; A character…Lily… ”
Ground Truth Resp A.
RRM-7B “ …both responses are within that limit …Assistant 2’s response is more detailed…more vivid and unique …  (Choose B)
JudgeLRM “ …In contrast, Assistant 2 provided a more detailed and vivid description, painting a picture of a character named Lily who is not only a leader but also a survivor, a fighter, and a free spirit …  (Choose B)
Rubric-RM
Rubric “1. The response uses strong imagery and creative language to create a vivid and unique character description. [Hard Rule] 2. The response is written in fewer than two paragraphs. [Hard Rule] 3. The response presents a character with distinctive and memorable traits that contribute to a clear and engaging portrayal. [Principle] 4. The response employs sensory details and evocative descriptions to enhance the reader’s mental image. [Principle] 5. The response demonstrates originality and creativity in language and characterization to avoid clichés and generic portrayals. [Principle] 6. The response maintains coherence and focus, ensuring all elements contribute to a unified and compelling character depiction. [Principle] 7. The response balances detail and conciseness, providing enough information to be vivid without unnecessary elaboration. [Principle] ”
Judge Identified Gatekeeper Criterion: The response is written in fewer than two paragraphs. …Response A successfully adheres to all criteria …Response B fails the paragraph count rule …While Response B offers more vivid and creative language …  (Choose A)
Table 5: Case study with error highlighting. Baselines favor the longer, imagery-rich response (B) and miss the explicit paragraph limit, while Rubric-RM first enforces hard rules and then scores principles.
We end up this section with concrete case studies on how Rubric-RM addresses challenging inputs and leads to better reward modeling. We present two instances drawn from RewardBench and FollowBench benchmarks, with baselines chosen from same-size generative reasoning reward models. Namely, we use JudgeLRM-7B and RRM-7B in Case 1 study, and two RM-R1-7B variants (DeepSeek-Dist, Qwen-2.5-Inst) for Case 2 study. The detailed results are shown in Table 5 and Table 6 respectively.
Case Study: Instruction adherence on FollowBench
Resp B (snippet) “‘…Would you consider direct air carbon capture as a significant financial venture? Yes …a recent report from BloombergNEF, published on May 16, 2024, titled …BNEF’s projection highlights several key economic implications: 1. …”
Ground Truth Resp B.
RM-R1-7B (DS) “ …looking at Chatbot A’s response. It starts by explaining DAC briefly and then moves into the study from Nature Energy in 2023 …Chatbot B’ …The quote is from BloombergNEF, but the publication date isn’t provided …  (Choose A)
Rubric-RM
Rubric “1. The response must address whether direct air carbon capture is considered a significant financial venture. [Principle] 2. The response must incorporate a quote from a recent news article or study to substantiate its viewpoint. [Hard Rule] 3. The response must concisely summarize the quoted source. [Hard Rule] 4. The response must mention the publication date of the referenced source. [Hard Rule] 5. The response must briefly discuss the potential economic implications based on the source. [Hard Rule] 6. The response must be written in a clear and understandable manner. [Principle] 7. The response must be well-organized and easy to follow. [Principle] 8. The response must be free of factual errors. [Principle] 9. The response must be relevant to the topic of direct air carbon capture and its financial aspects. [Principle] ”
Judge Identified Gatekeeper Criterion: The response must incorporate a quote from a recent news article or study to substantiate its viewpoint. …Response B is slightly better organized …for the economic implications …also provided a more recent source …  (Choose B)
Table 6: Case study 2 with error highlighting. Baselines overvalue academic-looking prose and misreading the date requirement, whereas Rubric-RM prioritizes hard rules and makes correct decision.
Case 1 (RewardBench–Chat Hard): instruction adherence vs. verbosity bias.
Both responses contain vivid descriptions, but the instruction explicitly requires fewer than two paragraphs. Baseline reasoning RMs ignore this hard requirement and choose the longer response, exhibiting a classic verbosity bias and instruction-violation blindness. In contrast, Rubric-RM first applies the Gatekeeper check (paragraph count), rejects the non-compliant candidate, and then scores principles (imagery/originality/focus), ultimately selecting the correct answer. This example highlights that long chain-of-thought does not guarantee correct constraint satisfaction, whereas a rubricjudge decomposition makes the failure explicit and avoids it.
Case 2 (FollowBench): verifiable recency and citation integrity.
This example is more challenging: both answers are longer and the quality gap is subtle. Baselines nevertheless produce factual mistakes about the evidence, e.g., asserting that the better response lacks a date/citation, despite it correctly providing a BloombergNEF quote with a May 16, 2024 publication date and concrete figures (\$387B cumulative investment). Our rubric-aware judge identifies recency and verifiability as hard requirements (quote, date, concise summary, and economic implications), and favors the response that meets them. This demonstrates Rubric-RM’s robustness to citation hallucinations and over-weighting of “academic-looking” prose that misled generative reasoning RMs.
Takeaways.
Across both cases, baseline reward models with long CoT capabilities still fail for two recurring reasons: (i) they overlook explicit hard rules (structural and evidentiary constraints), and (ii) they are vulnerable to hallucinated or weakly verifiable references. By contrast, Rubric-RM enforces gatekeepers before scoring higher-level qualities, yielding interpretable decisions and improved accuracy on difficult examples. We also observe that the gatekeeper stage reduces false positives from off-task/overlong content and elevates verifiability-aware judging in domains (e.g., science/finance) where recency and source integrity matter.

### v2

@@ 5.1 Datasets and Experiment Settings
Training data. We train both components of Rubric-RM: the rubric generator and the judge, on the curated OpenRubrics as presented in Sec. 4.2. Rubrics are produced with contrastive signals from chosen/rejected responses and filtered by preference–label consistency before use. Unless otherwise noted, we use the science-related slice of OpenRubrics to better match our domain study on HealthBench/medical evaluation.
@@
Backbone and variants. Both the rubric generator and the judge are fine-tuned from Qwen-3-8B (“Rubric-RM-8B”) unless specified. At inference time, Rubric-RM follows the two-stage process as detailed in Sec 4.3. We also report an ensemble variant, voting@5, which aggregates five independently sampled judge trajectories by majority vote.
@@
Baselines. We compare Rubric-RM against strong, same-scale white-box reward (judge) models: JudgeLRM-7B (chen2025judgelrm), RRM-7B (guo2025reward), and RM-R1-7B (chen2025rm). We also report larger RM-R1-14B (chen2025rm) and API judges for reference when available. To isolate the benefit of rubric-aware fine-tuning, we include naive Qwen-3-8B (Rubric+Judge) that directly prompts the base model to produce rubrics and then make judgments.
@@
Evaluation benchmarks and metrics. We evaluate Rubric-RM as a pairwise reward model on popular reward-modeling benchmarks: RewardBench (Chat, Chat-Hard) (rewardbench), RM-Bench (liu2025rmbench), PPE-IFEval (ppe), FollowBench (followbench), InfoBench (infobench), IFBench (ifbench), and RewardBench2 (Precise-IF, Focus) (malik2025rewardbench2). While FollowBench and InfoBench were originally designed to assess instruction-following capabilities of LLMs, we adapt them to pairwise evaluation settings by sampling two responses from the same model (Qwen-3-8B/14B), where one response adheres to all specified constraints and the other violates some of them. For the domain study we additionally report HealthBench/medical results. We follow each benchmark’s official splits and scoring rules, reporting accuracy, win-rate or other specific scores.
Due to space limits, additioanl implementation details are deferred to Appendix B.
RewardBench IF Evaluation Benchmarks RM-Bench RewardBench2 HelpSteer3 Avg.
@@
Qwen-3-8B (Rubric+Judge) 74.2 64.2 71.3 57.0 74.3 59.5 63.9 8.1 44.0 60.8 57.7
@@
Rubric-RM-4B 87.2 68.9 78.0 64.8 79.3 63.2 62.6 35.6 79.6 64.7 68.4
Rubric-RM-4B-voting@5 88.7 70.2 80.7 66.2 83.0 64.9 63.8 38.1 82.6 65.0 70.3
Rubric-RM-8B 88.2 74.1 76.1 67.0 80.8 65.4 65.7 34.4 82.2 67.0 70.1
Rubric-RM-8B-voting@5 89.9 75.4 81.5 70.8 83.8 67.1 67.0 40.0 86.5 67.5 73.0
5.2 Performance of Rubric-RM
We first validate the performance of Rubric-RM for reward modeling. For a more systematic evaluation, we test both 4B and 8B variants of Rubric-RM, which use Qwen3-4B/8B as backbones. Table 1 reports the results of Rubric-RM.
Outperforming Comparable Reward Models. Both Rubric-RM-4B and 8B surpass all existing 7B-scale white-box baselines (e.g., JudgeLRM, RRM, RM-R1). Notably, Rubric-RM-4B (68.4) outperforms the strongest 7B competitors (max 61.7), with Rubric-RM-8B further improving to 70.1. This confirms that rubric-aware training yields more reliable signals than generic preference learning, even with reduced parameter counts.
Majority Voting Further Enhances Performance. We also evaluate Rubric-RM-voting@5, which aggregates predictions via majority voting across five independent judge trajectories. Rubric-RM-4B-voting@5 reaches 70.3, and Rubric-RM-8B-voting@5 achieves the best overall average of 73.0, surpassing much larger models such as RM-R1-14B (71.7) and the Rubric+Judge API (71.3). These results highlight the stability benefits of rubric-based ensembles.
Effectiveness of Rubric-Aware Fine-Tuning. Directly using Qwen-3-8B to generate rubrics and judge yields only 57.7. In contrast, Rubric-RM achieves 70.1. This +12.4 gain validates our core contribution: high-quality rubrics derived via contrastive generation and consistency filtering are essential for effective reward modeling.
Strength on IF Benchmarks. Rubric-RM excels on benchmarks requiring fine-grained instruction adherence. On FollowBench and InfoBench, we achieve 81.5 and 83.8 respectively, substantially outperforming baselines like JudgeLRM and RRM. This demonstrates that rubrics capture nuanced constraints better than scalar reward models.
In the remaining experiments, we use Rubric-RM-8B as our reward model unless specified.
@@ 5.2 Performance of Rubric-RM
AutoIF (dong2024self) 56.9 47.1 67.0 57.6 57.2 80.6
UltraIF (an2025ultraif) 75.4 71.3 83.0 79.4 77.3 80.7
@@
+ DPO (via Rubric-RM) 78.2 73.9 84.5 81.2 79.5 83.0
Table 2: Comparison of trained policy models with different reward models on a format-based instruction-following benchmark (IFEval) and an open-ended benchmark (InfoBench). Baseline results are from viswanathan2025checklists. Results with underlines are reproduced by us using official checkpoints and evaluation scripts. Best scores are in bold.
@@
Vanilla SC Vanilla LC
@@
+ DPO (via Rubric-RM) 52.9 53.1 47.0 41.3 48.6
Table 3: Comparison of different alignment strategies applied to Qwen2.5-7B-Instruct on Arena-Hard and AlpacaEval. Baseline results are from viswanathan2025checklists. SC and LC stands for Style-controlled and Length-controlled. Best results are in bold.
Figure 4: Comparison of policy models on IFBench. Results of baselines except RLCF are from bhaskar2025language. We evaluate RLCF with its official checkpoint.
Method Creative Planning Math Info seeking Coding WB Score
Claude-3.5-Sonnet 55.6 55.6 50.2 55.5 56.5 54.7
GPT-4-turbo 58.7 56.2 51.0 57.2 55.1 55.2
GPT-4o-mini 60.1 58.2 54.0 57.4 57.2 57.1
Qwen2.5-7B-Instruct 50.1 51.8 47.1 50.7 45.0 48.7
+DRIFT 52.5 53.2 50.6 52.4 50.3 51.7
+DPO (via RLCF) 51.4 52.7 49.0 51.3 48.8 50.5
+DPO (via RLMT (PPO)) 52.1 52.6 45.2 51.4 48.3 49.7
+DPO (via Rubric-RM) 54.8 55.5 51.5 54.1 52.9 53.6
Table 4: Comparison of different alignment strategies applied to Qwen2.5-7B-Instruct on WildBench. Results are reported for task-specific scores and task macro WB score. Baseline results with "*" are from wang2025drift. We evaluate RLCF and RLMT with their official checkpoints. Best results are in bold.
5.3 Policy Models with Rubric-RM
@@ 5.3.1 Instruction-Following Evaluation
We further evaluate the effectiveness of using Rubric-RM as a reward model for policy optimization on instruction-following tasks, including IFEval, InfoBench, and IFBench. The results are shown in Table 2 and Figure 4.
Improved Performance on IFEval and InfoBench. Using Rubric-RM yields the best overall performance among open-source reward-model baselines: the resulting policy model reaches 79.5 on IFEval (vs. 76.0 with Skywork/ArmoRM) and 83.0 on InfoBench, approaching much larger commercial systems. These results indicate that rubric-based rewards provide more reliable optimization signals for constrained instruction following.
Clear Gains on Complex Instruction Following (IFBench). Figure 4 shows that the policy model optimized with Rubric-RM achieves 30.3 on IFBench, substantially higher than RLCF (28.2) and RLMT-based methods (22.4–22.8). Compared with both supervised fine-tuning variants and reinforcement learning baselines, Rubric-RM provides stronger inductive biases, enabling policies to better capture fine-grained instruction adherence.
Overall, these results demonstrate that Rubric-RM, when used as a reward model, provides a substantially stronger training signal that improves the instruction-following capability of learned policies.
5.3.2 General Alignment Evaluation
We evaluate policy model trained with Rubric-RM on alignment benchmarks Arena-Hard, AlpacaEval, and WildBench (Tables 3 and 4).
With DPO optimization, Rubric-RM achieves the best overall average score (48.6) among all open-source reward models. On Arena-Hard (style-controlled), it obtains 53.1, outperforming Skywork (50.3), Ultrafeedback (47.9), and RLCF (48.4). On AlpacaEval, it reaches 47.0, surpassing Skywork (44.8) and ArmoRM (37.6). On WildBench, aligning Qwen2.5-7B-Instruct with Rubric-RM yields the best macro WB score of 53.6, improving over the base model (48.7) by +4.9 and outperforming DPO with RLCF (50.5), RLMT (49.7), and DRN (51.7), with consistent gains across all task categories. These results show that rubric-based signals provide reliable gains across both vanilla and controlled settings.
@@ 5.4 Rubric-RM for BioMedical Domain
In this section we further study the effectiveness of Rubric-RM in more specialized biomedical domain, following arora2025healthbench. The Rubric and Judge are Qwen-3-8B backbones fine-tuned with OpenRubrics data from science-related domains.
Performance on HealthBench. Rubric-RM outperforms comparable-size reasoning reward models on HealthBench, achieving 68.3, exceeding RRM-7B (63.3) and RM-R1-7B variants (55.4/66.9), and approaching RM-R1-14B (69.9). Majority voting further boosts performance: Rubric-RM-voting@5 reaches 72.9, narrowing the gap to 14B reasoning models and API-based judges. A key finding is the importance of domain- specific rubric SFT. Directly prompting Qwen-3-8B with a Rubric+Judge pipeline yields only 51.8, whereas Rubric-RM improves performance by +16.5. This highlights our core contribution: contrastive, domain-tuned rubric training produces higher-precision, science-aware evaluation signals than on-the-fly rubric generation.
@@
Preference Optimization with Rubric-RM on HealthBench. We use Rubric-RM as the preference judge for DPO on HealthBench. Starting from Qwen-2.5-7B-Instruct (21.6), DPO with ArmoRM and RM-R1-7B improves performance to 22.5 and 22.7, respectively. In contrast, DPO with Rubric-RM achieves the best result at 23.8, yielding a consistent +1.1–1.3 gain over strong 7B reasoning rewards. These results demonstrate that rubric-aware, domain-specialized reward models translate directly into stronger biomedical policy learning, outperforming generative reasoning rewards.
Case Study on RewardBench Chat Hard
Prompt Describe a vivid and unique character, using strong imagery and creative language. Please answer in fewer than two paragraphs.
Resp B (snippet) “…a fierce and determined young woman …[paragraph break] …Which of the following is an example of a character …[paragraph break] A character …Lily …”
Label Resp A.
RRM-7B “…both responses are within that limit …Assistant 2’s response is more detailed …(Choose B)
JudgeLRM “…Assistant 2 provided a more vivid description …(Choose B)
Rubric-RM
Rubric “1. Describe a vivid and unique character. [Hard Rule] 2. Use strong imagery and creative language. [Hard Rule] 3. Be fewer than two paragraphs. [Hard Rule] 4. Employ sensory-rich language. [Principle] 5. Use original expressions. [Principle] 6. Maintain a consistent tone. [Principle] 7. Remain cohesive and focused. [Principle] 8. Avoid unnecessary elaboration. [Principle]”
Judge “The paragraph limit is the most verifiable constraint. Response B fails the paragraph rule, while Response A satisfies all hard rules and most principles …(Choose A)
Table 5: Case study with error highlighting. Baselines favor the longer, imagery-rich response and miss the explicit paragraph constraint, while Rubric-RM enforces hard rules before evaluating principles.
@@ 5.5 Efficiency Comparison
Table 6: Computing speed on 100 samples (vLLM).
Table 6 reports wall-clock time on 100 randomly sampled prompts from RewardBench2. Despite using two Qwen-3-8B components (rubric generator + judge), Rubric-RM runs in 130.77s, which is no slower than existing reasoning reward models, including RRM-7B (203.4s) and RM-R1-7B/14B (170.8–382.0s). It is substantially faster than 14B R1 variants and competitive with strong 7B reasoning baselines. Efficiency stems from our architecture: rather than long Chain-of-Thought decoding, we use short rubric generation followed by a lightweight judge. Moreover, rubrics are amortizable and can be cached for reuse across many examples, removing rubric generation cost during large-scale scoring. Although JudgeLRM achieves lower latency, it lacks the explicit, interpretable signals necessary for downstream policy optimization.
5.6 Case Studies
We conclude this section with a case study that illustrates how Rubric-RM handles challenging inputs and improves reward modeling. Table 5 shows the case study on RewardBench, where both responses are vivid, but the instruction requires fewer than two paragraphs. The baselines miss this hard constraint and incorrectly prefer the longer answer, exhibiting a verbosity bias and an instruction violation. In contrast, Rubric-RM first applies a simple structural check to reject the non-compliant candidate, then compares higher-level qualities (e.g., imagery, originality, focus), and selects the correct response. This example shows how long CoT can still overlook explicit constraints, while rubric-based decomposition makes such failures clear.